# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ III Architecture and Design Overview
The substrate to create GKP and magic states is a cluster state formed from two-mode squeezed light and passive beamsplitters, where each node is a photonic mode characterized by the quadratures $q$ and $p$, and features nearest-neighbor entanglement between the nodes. The time-multiplexed cluster state in our architecture is indexed by spatial and optical frequencies along with temporal modes (see Fig. 4 and later figures on cluster state engineering). To produce cat states our architecture uses PNR detection at specific optical frequencies and timesteps. Specifically, Schrödinger cat states are produced using the Photon-counting-Assisted Node- Teleportation Method (PhANTM) Eaton et al. (2022a). PhANTM is well-suited to our time-frequency-space multiplexed architecture, as individual nodes of a cluster state can be made non-Gaussian with photon subtraction, and the non-Gaussianity can be teleported through a cluster state with homodyne measurements. This allows us to repeat the PhANTM process over multiple time steps to increase the amplitude of the generated cats and make them suitable for high-probability GKP generation, in contrast to single-shot GKP generation methods without time or frequency multiplexing Bourassa et al. (2021); Takase et al. (2024).
@@
The cat states, which are embedded in the cluster state, are squeezed using teleportation-based squeezing gates Alexander et al. (2014) and then “bred” into either sensor states for qubit production or GKP states for magic state production Baragiola et al. (2019). Switching between the two output states is achieved by changing the required amplitude of the input cat states. Given that our protocols produce cat states with probabilistic squeezing and amplitude located at the nodes of a cluster state, we use an adaptive breeding protocol to produce GKP qubits or sensor states. In our adaptive protocol, cat states of various multiples of the desired grid state spacing are used including a single mode squeezed vacuum that can be substituted for lower quality cat states when desired (see Algorithm  V).
@@
These fundamental components can all be implemented in passive photonic integrated circuits, with squeezed light, sensor states, magic states, and the RHG lattice all produced on integrated chips. The chips integrate linear optics components, such as beamsplitters and phase shifters, and nonlinear components such as $\chi^{(3)}$ media, as shown in Fig. 3. A squeezing resonator is bidirectionally pumped by pumps centered at $\nu_{0}$ with $\nu_{p_{1}}=\nu_{0}+\Delta\nu/2$ and $\nu_{p_{2}}=\nu_{0}-\Delta\nu/2$, where $\Delta\nu$ is the free spectral range (FSR) of the resonator. This creates two squeezed combs each centered about the respective pump. Interfering the two squeezed combs at a balanced beamsplitter results in the formation of a 1D dual-rail quantum wire Alexander et al. (2016). Subsequently, homodyne measurement are performed locally within this cluster to reshape it in preparation to apply PhANTM or adaptive breeding at different time steps. If the local oscillator phase is aligned with the $q$ quadrature then the node is removed whereas if the phase is that of the $p$ quadrature, the quantum information is teleported to a neighboring node, with an additional $\pi/2$ rotation applied Menicucci et al. (2006). Each frequency mode needs to be separated from the optical frequency comb using frequency filters to then be detected. These filters are schematically shown as ring resonators, which are among the most compact and lowest-loss options for frequency separation Liu et al. (2021) negligibly impacting the pass-through frequency modes, and demonstrating very high transmission efficiencies of the filtered mode. The delay is used for time multiplexing before a final balanced beamsplitter is used to interfere frequency modes in (b) with the delayed modes in (a) to produce the final cluster state. With this resource state CVQC operations can be applied to generate cat states by performing PhANTM, squeezed cats using a squeezing gate, or GKP states by adaptive breeding.
@@
The architecture implements quantum error correction (QEC) by embedding the bred GKP sensor states from the photonic integrated circuit in Fig. 3 into the RHG cluster state Raussendorf et al. (2006), the MBQC analog of the surface code. Large scale entanglement is generated using a macronode based approach Tzitrin et al. (2021) where each qubit of the RHG lattice is replaced by four optical modes. Entanglement links in the RHG lattice are replaced by GKP Bell pairs and the final cluster state is then generated by interfering all modes of the macronode with passive linear optics.
Logical gates can be implemented on the RHG cluster state using lattice surgery Herr et al. (2018) and magic states. Magic states can be embedded in the macronode RHG lattice by replacing a GKP Bell pair with the result of interfering a magic state with squeezed vacuum Tzitrin et al. (2021). More detail on the lattice construction is given in Sec. VII.
IV Cat States on dual-rail quantum wire
Cat states are a fundamental resource in our CVQC architecture. They are used as a resource state to generate both sensor states via breeding and magic states via implementing CV-QEC on input cat states. Methods for creating cat states in CV photonics need at a minimum a combination of Gaussian and non-Gaussian resource states, or photon number detection along with Gaussian resource states. A method was proposed to create cat states by applying PNR on squeezed vacuum using machine learning to control some of the experimental parameters Anteneh et al. (2024). This method is efficient for state generation, requiring only a few iterations of photon detection, but requires high photon number resolution. Another proposal used high-photon-number Fock states (with $10$ or more photons) along with photon number resolution to create cat states deterministically Winnel et al. (2024). Creating high photon-number Fock states presents a high experimental challenge Uria et al. (2020). In PhANTM Eaton et al. (2022a), the method used in the present architecture, applications of probabilistic photon subtraction and teleportation build up non-Gaussianity over time. This requires a Gaussian cluster as a resource to repeatedly apply the method. One of the key advantages of PhANTM lies in its ability to generate sufficiently large cat states that can subsequently be bred while distributing the total number of photons subtracted over multiple repetitions of the operation.
@@ IV Cat States on dual-rail quantum wire
The application of PhANTM on the dual-rail wire in Fig. 4 IV corresponds to repeating the circuit in Fig. 5 in time. The rotation on the output mode can be manipulated through the phase of the beamsplitter and homodyne detection basis. The beamsplitter is defined by two parameters, $\theta$ and $\phi$, which are related to the reflectivity and transmitivity such that $r=e^{i\phi}\sin(\theta)$ and $t=\cos(\theta)$. We have chosen the phase $\phi=\pi/2$ in Fig. 5 such that a single application of the circuit at $\tau_{i}$ produces an output cat with the same orientation as the input cat, allowing repeated applications of PhANTM at $\tau_{i+1}$ and later timesteps. Note that there are two homodyne measurements applied in a unit round of PhANTM (one on $\nu_{a}$ and another on $\nu_{b}$) before the output mode is ready for the next round. Also note that given the beamsplitters and phases involved in the circuit in Fig. 5, the bases used for either mode teleportation or deletion differ from the usual $p$ or $q$ bases for a canonical cluster state  Raussendorf (2003). The effective $C_{Z}$ gate in Fig. 5 results from the application of two-mode squeezing, beamsplitters, and homodyne measurements as shown in Fig. 4. The modes that undergo operation $\mathcal{O}$ in Fig. 3 have multiple photon subtractions followed by homodyne detection applied to them, while the rest of the modes in IV have $p$ basis homodyne measurement applied to them. Active displacements are needed in between repeated applications of PhANTM due to random homodyne measurement results. The value of the displacement is the output of the homodyne detector fed directly back to a standard displacement circuit with a local oscillator for the next temporal mode Eaton (2022). The timescale of these active displacements must be shorter than the time delay between successive temporal modes in the cluster state.
@@
Monte Carlo simulation of several rounds of PhANTM were conducted using the unit cell shown in Fig. 5, sampling the number of photons detected in each photon subtraction. The parameters of the cat states that we aim to optimize are their size and squeezing, both of which are stochastic. The larger these parameters, the higher the effective squeezing of the resulting GKP states, leading to lower error rates for QEC codes. To facilitate analysis and comparison across various conditions, a single parameter is defined: the amplitude of the cat state with zero squeezing ($\alpha_{c}$, also see Sec. V). Fig. 6 shows the mean of $\alpha_{c}$ as a function of cluster squeezing for different number of PhANTM rounds (see also Sec. XI.3 in Appendix). In these simulations, eight sequential photon subtraction attempts were performed at each time step. The beamsplitter angles for each photon subtraction are chosen such that no more than $10$ photons are detected per PNR detector. The subtraction probability goes down for later subtraction events, as continued subtraction attempts de-amplify the quantum signal. We apply a reflectivity gradient (higher angle in later subtraction attempts) to ensure that the distribution of subtracted photons among the PNR detectors is roughly uniform (see Sec. XI.4 in Appendix). Increasing the cluster squeezing enhances the potential number of photons that can be subtracted in a single subtraction event, resulting in cats with larger amplitudes. Furthermore, increasing the number of PhANTM steps raises the total number of photons subtracted across all cumulative PhANTM events, leading to larger amplitudes for cat states. The additional squeezing value also depends on the reflection and transmission coefficients of each beamsplitter utilized in the photon subtraction process.
@@ V Generation of GKP States
The pseudoalgorithm for adaptive breeding is shown in Algorithm V. The algorithm takes a set of cat states embedded in a cluster state. In order to breed GKP or sensor states with the correct spacing, the cat amplitudes must be adjusted using a squeezing operator. Once the amplitudes are adjusted, the cats are simply measured in the $p$ basis. Because the cats are embedded in a cluster state, they have effective $C_{Z}$ operations applied between them, and squeezing and breeding is thus implemented via measurements only Eaton et al. (2022a).
@@
Input desired cat $\alpha_{b}$ in Weigand and Terhal (2018) $\alpha_{b}=\xi 2^{(M-3)/2}$
@@
         teleport with $p$
@@
         teleport with $p$
@@
         measure with $q$
@@
The adaptive breeding protocol takes several inputs: the number of desired breeding rounds, $M$; the squeezing present in the cluster state, $r_{cluster}$; a number of input cats which depends on the number of breeding rounds, $2^{M}$; the input cats’ unsqueezed amplitudes $\alpha$, their present squeezing level after PhANTM, $r_{cat}$, and their parity $P$; the desired GKP or sensor state spacing in phase space, $\xi$; the required cat state amplitude for breeding, $\alpha_{b}$, which depends on the spacing as $\alpha_{b}=\xi 2^{(M-3)/2}$ Weigand and Terhal (2018); and the lower bound on cat squeezing that will determine its suitability for breeding, $r_{lb}$. All measurements are performed in either the $p$ or $q$ basis depending on whether an input cat is used in breeding or replaced with a squeezed vacuum.
@@
As a first requirement for the adaptive breeding to proceed, the amplitude $\alpha$ of the cats output from PhANTM must be known. The output amplitude and squeezing, $r_{cat}$ depends on the photon subtraction vector obtained from PhANTM. The parity $P$ of each cat also depends on the total number of photons subtracted during PhANTM. As each cat can only breed with another cat of the same parity, small displacements are used to adjust the odd cat parities to even ones. This displacement causes an effective, known displacement in the first breeding round output, which can be kept track of but does not have to be actively fixed until breeding is complete. In general, the cat amplitudes are adjusted via squeezing operations of magnitude $s_{\alpha_{b}}$, which depends on the desired spacing $\xi$ and the number of breeding steps $M$.
@@
The extra teleportation step in the breeding protocol allows us to introduce momentum squeezed states at $r_{cluster}$ squeezing in two scenarios. If the cat amplitude produced by PhANTM is too small to be used in GKP generation such that no amount squeezing smaller than $r_{cluster}$ can be used to adjust its amplitude to $\alpha_{b}$, it can be replaced by a momentum squeezed state before breeding. Adding a squeezed vacuum state in place of a cat increases the squeezing in one quadrature at the expense of the squeezing in the other quadrature. In the second scenario, arbitrarily large cat states can be produced in PhANTM that are too big to be squeezed to amplitude $\alpha_{b}$ needed in breeding because the squeezing required to do so would exceed $r_{cluster}$. These cats can be squeezed to an even multiple of the required amplitude for the breeding protocol to proceed ($2\alpha_{b}$ is used in Algorithm V). If the squeezing of these large cats, $s_{2\alpha_{b}}$, is smaller than the desired lower bound, $r_{lb}$, then these cats can once again be replaced by momentum squeezed states (and likewise if the required squeezing would exceed $r_{cluster}$). The squeezing of cats or cluster is the dominant term that determines the effective squeezing of the GKP states in one quadrature, while the number of breeding rounds is the dominant term for effective squeezing in the other direction. Note that using a cat that is an even multiple of the required cat squeezing also has the effect of increasing the effective squeezing in one of the quadratures, giving us flexibility in trading off the squeezing in the GKP quadratures by combining mixtures of cats with different amplitudes and momentum squeezed states.
@@
Fig. 7 illustrates the cluster engineering method required to apply adaptive breeding in Algorithm V to cat states generated on different dual rails of the cluster state (see also Appendix Sec. XI.5). At stage II, the modes arrangement is modified compared with Fig. 4. Specifically, prior to applying any breeding operation, distinct frequency modes need to be connected via an effective $C_{Z}$ interaction. This connection is shown in stage III. Breeding of cat states in mode $4$ and $8$ is applied in stage IV at the time step $\tau_{i+2}$. Prior to adaptive breeding, squeezing gates are employed to ensure cat states sizes match $\alpha_{b}$ or an even multiple of $\alpha_{b}$ in the pseudoalgorithm above. This operation requires one single time step and consumes two spatial modes of the same frequency as it is represented between time step $\tau_{i+1}$ and $\tau_{i+2}$ in Fig. 7. The squeezing parameter of the gate is set by the homodyne measurement angles $\theta_{a}$ and $\theta_{b}$ Alexander et al. (2014). More details are discussed in Sec. XI.5 of the Appendix. Note that we can swap between sensor states and GKP qubits by changing the amplitudes to which the cats are squeezed.
@@
The outputs of the previously described PhANTM simulations are used as input to an adaptive breeding simulation. A Monte Carlo approach is used to obtain statistics about the sensor states obtained from adaptive breeding, with 1000 samples drawn from a set of input cats that were produced with an array of settings for starting cluster squeezing and the number of photon subtraction attempts. The simulation makes use of Qutip Johansson et al. (2013) for density matrix calculations and Strawberry Fields Killoran et al. (2019) for homodyne detection simulation. The average sensor state squeezing along with their standard deviations are shown alongside FT threshold simulations in Sec. VIII.
VI Generation of magic States

### v3

@@ 3 Architecture and Design Overview
The substrate to create GKP and magic states is a cluster state formed from two-mode squeezed light and passive beamsplitters, where each node is a photonic mode characterized by the quadratures $q$ and $p$, and features nearest-neighbor entanglement between the nodes. The time-multiplexed cluster state in our architecture is indexed by spatial and optical frequencies along with temporal modes (see Fig. 4 and later figures on cluster state engineering). To produce cat states our architecture uses PNR detection at specific optical frequencies and timesteps. Specifically, Schrödinger cat states are produced using the Photon-counting-Assisted Node- Teleportation Method (PhANTM) [55]. PhANTM is well-suited to our time-frequency-space multiplexed architecture, as individual nodes of a cluster state can be made non-Gaussian with photon subtraction, and the non-Gaussianity can be teleported through a cluster state with homodyne measurements. This allows us to repeat the PhANTM process over multiple time steps to increase the amplitude of the generated cats and make them suitable for high-probability GKP generation, in contrast to single-shot GKP generation methods without time or frequency multiplexing [26, 57].
@@
The cat states, which are embedded in the cluster state, are squeezed using teleportation-based squeezing gates [58] and then “bred” into either sensor states for qubit production or GKP states for magic state production [25]. Switching between the two output states is achieved by changing the required amplitude of the input cat states. Given that our protocols produce cat states with probabilistic squeezing and amplitude located at the nodes of a cluster state, we use an adaptive breeding protocol to produce GKP qubits or sensor states. In our adaptive protocol, cat states of various multiples of the desired grid state spacing are used including a single mode squeezed vacuum that can be substituted for lower quality cat states when desired (see Algorithm  5).
@@
These fundamental components can all be implemented in passive photonic integrated circuits, with squeezed light, sensor states, magic states, and the RHG lattice all produced on integrated chips. The chips integrate linear optics components, such as beamsplitters and phase shifters, and nonlinear components such as $\chi^{(3)}$ media, as shown in Fig. 3. A squeezing resonator is bidirectionally pumped by pumps centered at $\nu_{0}$ with $\nu_{p_{1}}=\nu_{0}+\Delta\nu/2$ and $\nu_{p_{2}}=\nu_{0}-\Delta\nu/2$, where $\Delta\nu$ is the free spectral range (FSR) of the resonator. This creates two squeezed combs each centered about the respective pump. Interfering the two squeezed combs at a balanced beamsplitter results in the formation of a 1D dual-rail quantum wire [59]. Subsequently, homodyne measurement are performed locally within this cluster to reshape it in preparation to apply PhANTM or adaptive breeding at different time steps. If the local oscillator phase is aligned with the $q$ quadrature then the node is removed whereas if the phase is that of the $p$ quadrature, the quantum information is teleported to a neighboring node, with an additional $\pi/2$ rotation applied [19]. Each frequency mode needs to be separated from the optical frequency comb using frequency filters to then be detected. These filters are schematically shown as ring resonators, which are among the most compact and lowest-loss options for frequency separation [60] negligibly impacting the pass-through frequency modes, and demonstrating very high transmission efficiencies of the filtered mode. The delay is used for time multiplexing before a final balanced beamsplitter is used to interfere frequency modes in (b) with the delayed modes in (a) to produce the final cluster state. With this resource state CVQC operations can be applied to generate cat states by performing PhANTM, squeezed cats using a squeezing gate, or GKP states by adaptive breeding.
@@
The architecture implements quantum error correction (QEC) by embedding the bred GKP sensor states from the photonic integrated circuit in Fig. 3 into the RHG cluster state [56], the MBQC analog of the surface code. Large scale entanglement is generated using a macronode based approach [27] where each qubit of the RHG lattice is replaced by four optical modes. Entanglement links in the RHG lattice are replaced by GKP Bell pairs and the final cluster state is then generated by interfering all modes of the macronode with passive linear optics.
Logical gates can be implemented on the RHG cluster state using lattice surgery [61] and magic states. Magic states can be embedded in the macronode RHG lattice by replacing a GKP Bell pair with the result of interfering a magic state with squeezed vacuum [27]. More detail on the lattice construction is given in Sec. 7.
4 Cat States on Dual-Rail Quantum Wire
Cat states are a fundamental resource in our CVQC architecture. They are used as a resource state to generate both sensor states via breeding and magic states via implementing CV-QEC on input cat states. Methods for creating cat states in CV photonics need at a minimum a combination of Gaussian and non-Gaussian resource states, or photon number detection along with Gaussian resource states. A method was proposed to create cat states by applying PNR on squeezed vacuum using machine learning to control some of the experimental parameters [62]. This method is efficient for state generation, requiring only a few iterations of photon detection, but requires high photon number resolution. Another proposal used high-photon-number Fock states (with $10$ or more photons) along with photon number resolution to create cat states deterministically [63]. Creating high photon-number Fock states presents a high experimental challenge [64]. In PhANTM [55], the method used in the present architecture, applications of probabilistic photon subtraction and teleportation build up non-Gaussianity over time. This requires a Gaussian cluster as a resource to repeatedly apply the method. One of the key advantages of PhANTM lies in its ability to generate sufficiently large cat states that can subsequently be bred while distributing the total number of photons subtracted over multiple repetitions of the operation.
@@ 4 Cat States on Dual-Rail Quantum Wire
The application of PhANTM on the dual-rail wire in Fig. 4 IV corresponds to repeating the circuit in Fig. 5 in time. The rotation on the output mode can be manipulated through the phase of the beamsplitter and homodyne detection basis. The beamsplitter is defined by two parameters, $\theta$ and $\phi$, which are related to the reflectivity and transmitivity such that $r=e^{i\phi}\sin(\theta)$ and $t=\cos(\theta)$. We have chosen the phase $\phi=\pi/2$ in Fig. 5 such that a single application of the circuit at $\tau_{i}$ produces an output cat with the same orientation as the input cat, allowing repeated applications of PhANTM at $\tau_{i+1}$ and later timesteps. Note that there are two homodyne measurements applied in a unit round of PhANTM (one on $\nu_{a}$ and another on $\nu_{b}$) before the output mode is ready for the next round. Also note that given the beamsplitters and phases involved in the circuit in Fig. 5, the bases used for either mode teleportation or deletion differ from the usual $p$ or $q$ bases for a canonical cluster state  [65]. The effective $C_{Z}$ gate in Fig. 5 results from the application of two-mode squeezing, beamsplitters, and homodyne measurements as shown in Fig. 4. The modes that undergo operation $\mathcal{O}$ in Fig. 3 have multiple photon subtractions followed by homodyne detection applied to them, while the rest of the modes in IV have $p$ basis homodyne measurement applied to them. Active displacements are needed in between repeated applications of PhANTM due to random homodyne measurement results. The value of the displacement is the output of the homodyne detector fed directly back to a standard displacement circuit with a local oscillator for the next temporal mode [66]. The timescale of these active displacements must be shorter than the time delay between successive temporal modes in the cluster state.
@@
Monte Carlo simulation of several rounds of PhANTM were conducted using the unit cell shown in Fig. 5, sampling the number of photons detected in each photon subtraction. The parameters of the cat states that we aim to optimize are their size and squeezing, both of which are stochastic. The larger these parameters, the higher the effective squeezing of the resulting GKP states, leading to lower error rates for QEC codes. To facilitate analysis and comparison across various conditions, a single parameter is defined: the amplitude of the cat state with zero squeezing ($\alpha_{c}$, also see Sec. 5). Fig. 6 shows the mean of $\alpha_{c}$ as a function of cluster squeezing for different number of PhANTM rounds assuming $r_{2}=r_{3}$ (see also Sec. 11.3 in Appendix). In these simulations, eight sequential photon subtraction attempts were performed at each time step. The beamsplitter angles for each photon subtraction are chosen such that no more than $10$ photons are detected per PNR detector. The subtraction probability goes down for later subtraction events, as continued subtraction attempts de-amplify the quantum signal. We apply a reflectivity gradient (higher angle in later subtraction attempts) to ensure that the distribution of subtracted photons among the PNR detectors is roughly uniform (see Sec. 11.4 in Appendix). Increasing the cluster squeezing enhances the potential number of photons that can be subtracted in a single subtraction event, resulting in cats with larger amplitudes. Furthermore, increasing the number of PhANTM steps raises the total number of photons subtracted across all cumulative PhANTM events, leading to larger amplitudes for cat states. The additional squeezing value also depends on the reflection and transmission coefficients of each beamsplitter utilized in the photon subtraction process.
@@ 5 Generation of GKP States
The pseudoalgorithm for adaptive breeding is shown in Algorithm 5. The algorithm takes a set of cat states embedded in a cluster state. In order to breed GKP or sensor states with the correct spacing, the cat amplitudes must be adjusted using a squeezing operator. Once the amplitudes are adjusted, the cats are simply measured in the $p$ basis. Because the cats are embedded in a cluster state with effective $C_{Z}$ operations applied between the cluster modes, squeezing and breeding is implemented via measurements only [55].
@@
Input desired cat $\alpha_{b}$ in [54] $\alpha_{b}=\xi 2^{(M-3)/2}$
@@
         teleport to next time step
@@
         teleport to next time step
@@
         replace with momentum squeezed state
@@
The adaptive breeding protocol takes several inputs: the number of desired breeding rounds, $M$; the squeezing present in the cluster state, $r_{cluster}$; a number of input cats which depends on the number of breeding rounds, $2^{M}$; the input cats’ unsqueezed amplitudes $\alpha$, their present squeezing level after PhANTM, $r_{cat}$, and their parity $P$; the desired GKP or sensor state spacing in phase space, $\xi$; the required cat state amplitude for breeding, $\alpha_{b}$, which depends on the spacing as $\alpha_{b}=\xi 2^{(M-3)/2}$ [54]; and the lower bound on cat squeezing that will determine its suitability for breeding, $r_{lb}$. All homodyne measurements are performed in the appropriate basis depending on whether an input cat is used in breeding or replaced with a squeezed vacuum.
@@
As a first requirement for the adaptive breeding to proceed, the amplitude $\alpha$ of the cats output from PhANTM must be known. The output amplitude and squeezing, $r_{cat}$ depends on the photon subtraction vector obtained from PhANTM. The parity $P$ of each cat also depends on the total number of photons subtracted during PhANTM. As each cat can only breed with another cat of the same parity, small displacements can be used to adjust the odd cat parities to even ones. This displacement causes an effective, known displacement in the first breeding round output, which can be kept track of but does not have to be actively fixed until breeding is complete. In general, the cat amplitudes are adjusted via squeezing operations of magnitude $s_{\alpha_{b}}$, which depends on the desired spacing $\xi$ and the number of breeding steps $M$.
@@
The extra teleportation step in the breeding protocol allows us to introduce momentum squeezed states at $r_{cluster}$ squeezing in two scenarios. If the cat amplitude produced by PhANTM is too small to be used in GKP generation such that no amount squeezing smaller than $r_{cluster}$ can be used to adjust its amplitude to $\alpha_{b}$, it can be replaced by a momentum squeezed state before breeding. Adding a squeezed vacuum state in place of a cat increases the squeezing in one quadrature at the expense of the squeezing in the other quadrature. In the second scenario, arbitrarily large cat states can be produced in PhANTM that are too big to be squeezed to amplitude $\alpha_{b}$ needed in breeding because the squeezing required to do so would exceed $r_{cluster}$. These cats can be squeezed to an even multiple of the required amplitude for the breeding protocol to proceed ($2\alpha_{b}$ is used in Algorithm 5). If the squeezing of these large cats, $s_{2\alpha_{b}}$, is smaller than the desired lower bound, $r_{lb}$, then these cats can once again be replaced by momentum squeezed states (and likewise if the required squeezing would exceed $r_{cluster}$). The squeezing of cats or cluster is the dominant term that determines the effective squeezing of the GKP states in one quadrature, while the number of breeding rounds is the dominant term for effective squeezing in the other direction. Note that using a cat that is an even multiple of the required cat squeezing also has the effect of increasing the effective squeezing in one of the quadratures, giving us flexibility in trading off the squeezing in the GKP quadratures by combining mixtures of cats with different amplitudes and momentum squeezed states.
@@
Fig. 7 illustrates the cluster engineering method required to apply adaptive breeding in Algorithm 5 to cat states generated on different dual rails of the cluster state (see also Appendix Sec. 11.5). At stage II, the mode arrangement is modified compared with Fig. 4. Specifically, prior to applying any breeding operation, distinct frequency modes need to be connected via an effective $C_{Z}$ interaction. This connection is shown in stage III. Breeding of cat states in mode $4$ and $8$ is applied in stage IV at the time step $\tau_{i+2}$. Prior to adaptive breeding, squeezing gates are employed to ensure that the size of each cat state matches $\alpha_{b}$ or an even multiple of $\alpha_{b}$ in the pseudoalgorithm above. This operation requires one single time step and consumes two spatial modes of the same frequency as it is represented between time step $\tau_{i+1}$ and $\tau_{i+2}$ in Fig. 7. The squeezing parameter of the gate is set by the homodyne measurement angles $\theta_{a}$ and $\theta_{b}$ [58]. More details are discussed in Sec. 11.5 of the Appendix. Note that we can swap between sensor states and GKP qubits by changing the amplitudes to which the cats are squeezed.
Notice that applying the squeezing operation or replacing cat states with momentum squeezed states does not require any active switching of the quantum states. Replacement of cat states is accomplished by setting $\theta_{a}=\pi/2$ and $\theta_{b}=0$ in homodyne measurements on the cluster (see Fig. 7). This causes the cluster mode in the next time step to be a momentum squeezed state, which is subsequently used for breeding, removing the need for active switching with an extra teleportation step.
@@
To quantitatively model the adaptive breeding step, the outputs of the previously described PhANTM simulations are used as input to an adaptive breeding simulation. A Monte Carlo approach is used to obtain statistics about the sensor states obtained from adaptive breeding, with 1000 samples drawn from a set of input cats that were produced with an array of settings for starting cluster squeezing and the number of photon subtraction attempts. The simulation makes use of Qutip [69] for density matrix calculations and Strawberry Fields [70] for homodyne detection simulation. The average sensor state squeezing along with their standard deviations are shown alongside FT threshold simulations in Sec. 8.
6 Generation of Magic States